how many electrons in an atom of Sb have ml=+3? ml=-2? and n=5?

how do you find these numbers?
i understand how to find l numbers.
Answer #1

We have 4 quantum numbers.

1. n = Principle quantum number = Which determines the shell and its value can be 1,2,3,4,....

2. l = Azimuthal quantum number = which determines the angular momentum and shape of orbital. Its value = n-1

For s orbital, l = 0

For p -orbital, l = 1

For d-orbital, l = 2 ... and so on

3. ml = Magnetic quantum number = determines the orientation and its value = -l, -l+1, -l +2......0.....l-2,l-1,l

4. ms = spin quantum number = determines electron spin and its allowed values are +1/2 and -1/2

Each orbital can have maximum of 2 electrons and number of electrons = 2n2

(a) When ml = +3

l can be maximum 3

Number of electrons = 2n2 = 2(3)2 = 18 and in specific ml = +3, maximum electrons = 2

(b) When ml = -2, l = 2

and maximum number of electrons 2(2)2 = 8 and in specific ml = -2, maximum electrons = 2

(c) n = 5

maximum number of electrons = 2(5)2 = 50